Please solve it by hand or typing, don't use EXCEL.

1. Suppose you have the expectations of the HPR on stock market as follows: State of Economy Recession Normal growth Expansion Probability 0.2 0.7 0.1 HPR -12% 8% 16% A. Compute the mean of HPR on stocks. (3pts) B. Compute the standard deviation of the HPR on stocks. (5pts)
Answer #1

a.Mean=Respective probability*Respective HPR

=(0.2*-12)+(0.7*8)+(0.1*16)=4.8%

b.

probability HPR probability*(HPR-Mean)^2
0.2 -12 0.2*(-12-4.8)^2=56.448
0.7 8 0.7*(8-4.8)^2=7.168
0.1 16 0.1*(16-4.8)^2=12.544
Total=76.16%

Standard deviation=[Total probability*(HPR-Mean)^2/Total probability]^(1/2)

=8.73%(Approx).